# Dark lens candidates from Gaia Data Release 3

### v1

@@ 5 Estimating lens parameters of the candidate dark events
We used the same approach as in Wyrzykowski et al. (2016), Mróz & Wyrzykowski (2021) and Kruszyńska et al. (2022) to estimate the mass and distance to the lens. We dubbed it the DarkLensCode [^2] https://github.com/BHTOM-Team/DarkLensCode and we explained this method in greater detail in Howil et al. (2024, in prep.). The DarkLensCode was used to find the posterior distribution of lens distance and lens mass, using PDFs of the photometric model parameters, and the Galactic model. The final estimates are the median values of obtained mass and distance PDFs. Here we will focus on the presentation and of resulting mass and distance estimates. We present the results in Table 1.
If we found more than one solution passed the criteria outlined in Section 4, we combined the resulting posterior distribution into one, randomly selecting 50 000 points for each solution. We used three mass functions as lens mass priors: Kroupa (2001) function that describes stars:
@@
and a $f(M)\sim M^{-1}$ corresponding to applying no prior on the lens function. The reported values of the lens mass and distance are median values of the posterior distribution, while their uncertainty is represented by the 16-th and 86-th quantiles.
We noticed that the mass function greatly affects the lens mass estimate. Using the Kroupa (2001) mass function results in lighter lenses at greater distances, and in turn more likely MS stars. In contrast, the Mróz et al. (2021) mass function produced more massive lenses at closer distances. This is because the Kroupa (2001) mass function is steeper and has a lower chance of producing massive lenses.
We also calculated the probability that the lens is dark and not an MS star. We compared the brightness of the blend from the microlensing model and compared it to the brightness of an MS star at an estimated distance from the lens, including ($P_{A}$) and excluding ($P_{0}$) extinction in that direction. The extinction $A_{G}$ used for the calculation of the observed magnitude of an MS was taken, when possible, from Gaia DR3 (ag_phot value). Otherwise, we calculated $A_{G}$ using the reddening maps of Schlafly & Finkbeiner (2011) and a relation of $E(B-V)$ with $A_{G}$ from Wang & Chen (2019). We compared the $A_{G}$ value to extinction in $u^{\prime}$, $g^{\prime}$, $r^{\prime}$, and $i^{\prime}$ filters from Schlafly & Finkbeiner (2011) to confirm its validity. When we included extinction, it resulted in a lower probability for a dark lens. All input parameters are available in machine-readable form in an online archive attached to this paper.
@@ 6 Discussion and conclusions
We found four lenses for which the probability $P_{A}$ for the dark lens scenario exceeded 80% when we looked only at the Kroupa (2001) mass function (GaiaDR3-ULENS-069, GaiaDR3-ULENS-088, GaiaDR3-ULENS-343, and GaiaDR3-ULENS-353) and included extinction. We can expand this list by six additional events when demanding that probability only the probability $P_{0}$ is larger than 80% (GaiaDR3-ULENS- 155, GaiaDR3-ULENS-212, GaiaDR3-ULENS-259, GaiaDR3-ULENS-270, GaiaDR3-ULENS-363). The probability $P_{0}$ was calculated without including the extinction. All but one of these events have Galactic coordinates located near the Galactic Center. The estimated distance for all but one of these nine objects would suggest that they are located in the Galactic disc, rather than the Galactic bulge. This would mean, that the extinction should be smaller than the value used in this work. This in turn means, that it’s less likely that the object would be an MS star, since the true brightness of an MS star $G_{\mathrm{MS,true}}$ should lie between two marginal values calculated in this paper, namely $G_{\mathrm{MS,0}}<G_{\mathrm{MS,true}}<G_{\mathrm{MS,A}}$. Additionally, these eight lenses would belong to the Galactic disc population and five of them are located in the nearest 2.5 kpc radius. Only one lens, GaiaDR3-ULENS-212, seems to be located near the Galactic bulge and would belong to that population.
If we assumed that all of those nine events are dark lenses, and instead follow a Mróz et al. (2021) mass function, we would end up with the majority of events with masses in the range of white dwarfs and neutron stars. Two objects have slightly larger masses: GaiaDR3-ULENS-069 could be consistent with a mass-gap object, and GaiaDR3-ULENS-259 could be a mass-gap object or a light BH. GaiaDR3-ULENS-259 has also the largest uncertainty for its mass, the interval.
@@
In the case of events where Gaia reported a proper motion, we decided to calculate the transverse velocity. We present results for selected nine events in Figure 3. The more massive objects seem to have lower velocities than the less massive ones, but the error bars are large and it is hard to draw any conclusive statements. Moreover, here we are most likely getting an estimate of a velocity coming from a Galactic prior. More accurate estimates will be possible, once we obtain an astrometric time series for these events.
Here we also measure only the most likely mass of candidate events. We cannot however confirm their nature as WDs, NS or BHs until we perform additional observations in other ranges of the electromagnetic radiation, especially X-rays and UV. They could be some unusual objects, such as quark stars, or products of primordial black hole (PBH) mergers with other objects. Abramowicz et al. (2018) and Abramowicz et al. (2022) show a mechanism, that could produce a low-mass BH from a moon-mass PBH ($10^{25}\,\mathrm{g}>M_{\mathrm{PBH}}>10^{17}\,\mathrm{g}$) collision with a NS. Such low-mass BH would be in the mass range of a NS.
It is worth noting, that these are only candidates for dark lenses, and their mass measurement will be possible only when we will include astrometric time series. This data will be available only with Gaia DR4, no sooner than the end of 2025. This is an exciting prospect, as Gaia may allow us to observe previously unseen stellar populations. This becomes even more promising with the approaching start of the Vera C. Rubin Observatory and its Legacy Survey of Space and Time (Ivezić et al., 2019), as well as the launch of the Roman mission (Akeson et al., 2019; Spergel et al., 2015). Rubin will allow us to select long-duration microlensing events from the entire Galactic plane, while Roman can provide us with high-cadence astrometric and photometric observations in the Galactic bulge.
@@ Acknowledgements.
KK would like to thank Monika Sitek, Drs. Etienne Bachelet, Mariusz Gromadzki, Przemek Mróz, Radek Poleski, Milena Ratajczak, Rachel Street, Paweł Zieliński, Przemysław Mikołajczyk, as well as Profs. Michał Bejger, Wojciech Hellwing, Katarzyna Małek, and Łukasz Stawarz. This work was supported from the Polish NCN grants: Harmonia No. 2018/30/M/ST9/00311, Daina No. 2017/27/L/ST9/03221, and NCBiR grant within POWER program nr POWR.03.02.00-00-l001/16-00. LW acknowledges MNiSW grant DIR/WK/2018/12 and funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No. 101004719 (OPTICON-RadioNET Pilot, ORP). This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement. We acknowledge ESA Gaia, DPAC and the Photometric Science Alerts Team (http://gsaweb.ast.cam.ac.uk/alerts).
@@ Appendix A Additional tables
In this appendix, we provide three tables:
@@
  • Table 3, which is the list of data sets used to obtain the final models for each of the 35 analysed events;
  • Table 4, which is the list of parameters of best-fitting solutions of the 35 analysed events. In this paper, the baseline magnitude and blending parameter are only provided for the $G$-band. A full, machine-readable version of this table, with all parameters, is available online.
Table 2: Table of the results of cross-match between 204 analysed Gaia DR3 events with other surveys.

### v2

@@ 6 Estimating lens parameters of the candidate dark events
We used the same approach as in Wyrzykowski et al. (2016), Mróz & Wyrzykowski (2021) and Kruszyńska et al. (2022) to estimate the mass and distance to the lens. We dubbed it the DarkLensCode [^8] https://github.com/BHTOM-Team/DarkLensCode. and we explained this method in greater detail in Howil et al. (2024). The DarkLensCode was used to find the posterior distribution of lens distance and lens mass, using PDFs of the photometric model parameters, and the Galactic model. The final estimates are the median values of obtained mass and distance PDFs. Here we will focus on the presentation and the resulting mass and distance estimates. We present the results in Tables LABEL:tab:lens_mass and, Table LABEL:tab:lens_mass_edge.
If we found more than one solution passed the criteria outlined in Section 4, we analysed them separately providing mass estimates for each solution.
The extinction $A_{G}$ was calculated following a method similar to one outlined in Fukui et al. (2019), but we used Gaia DR3 data instead. We selected all sources within a 30’ radius with a Renormalised Unit Weight Error parameter smaller than 1.4, parallax uncertainty smaller than 20% of the measured value, and available GSP-Phot solutions (Andrae et al., 2023). Then we calculated the mean and standard deviation of the ag_gspphot in 50 pc bins and fitted a fourth-order polynomial. We used the fitted polynomial as a function of extinction depending on the distance towards the lens or the source. If the distance to the lens or source was larger than 8 kpc, we used the calculated extinction value $A_{G}$ at 8kpc.
We did not know the distance to the source, so we assumed different maximum and minimum ranges. For events located towards the bulge, we initially assumed that the distance can be between 1 kpc and 12 kpc. For events located towards the disc, we chose the distance between 0.1 kpc and 8 kpc (GaiaDR3-ULENS-118) or 10 kpc (GaiaDR3-ULENS-259). When available, we used $\theta_{*}$ derived in Section 5. First, we randomly selected a distance to the source from the range described above and calculated the source radius $R_{\mathrm{S}}$. In the next step, we found the absolute magnitude of the source star. The procedure depended on the position of the source star in the CMD.
If the source star position on the CMD was located in the main sequence, we assumed the star was a dwarf. Using the tables from Pecaut & Mamajek (2013) provided on the author’s website [^9] https://www.pas.rochester.edu/~emamajek/ we found the corresponding value of absolute magnitude in $G$ band.
If the source star position on the CMD was located in the red giant clump, we assumed the star was a giant. For the absolute magnitude determination, we followed the information contained in van Belle et al. (2021). First, we found the $(V_{\mathrm{0}}-K_{\mathrm{0}})$ value corresponding to the radius based on the inverted relation presented in the paper with coefficients coming from Table 16. Then we used Equation 4 from van Belle et al. (2021) to find the effective temperature $T_{\mathrm{eff}}$ of the source star. We used the well-known relation $L_{\mathrm{bol}}=4\pi R^{2}_{\mathrm{S}}\sigma_{\mathrm{SB}}T^{4}_{\mathrm{eff}}$ to find the bolometric luminosity of the source star. To find the bolometric correction $BC_{\mathrm{G}}$ in $G$ band, we followed the recipe provided in Manteiga et al. (2018), Chapter 8.3.3. Finally, we could derive the absolute magnitude in $G$ following Equation 8.6 from Manteiga et al. (2018).
We then found the extinction $A_{G}$ and the observed magnitude of the source star at a selected distance. If the absolute value of the difference between the calculated source magnitude and the observed source magnitude from the microlensing model was smaller than the sum of the source magnitude uncertainty from the model and the derived source magnitude uncertainty, we accepted that source distance value.
We used three mass functions as lens mass priors: Kroupa (2001) function that describes stars:
@@
and a $f(M)\sim M^{-1}$ corresponding to applying no prior on the lens function. The reported values of the lens mass and distance are median values of the posterior distribution, while their uncertainty is represented by the 16-th and 84-th quantiles.
We noticed that the mass function greatly affects the lens mass estimate. Using the Kroupa (2001) mass function results in lighter lenses at greater distances, and in turn more likely MS stars. In contrast, the Mróz et al. (2021) mass function produced more massive lenses at closer distances. This is because the Kroupa (2001) mass function is steeper and less likely to produce massive lenses.
We compared the brightness of the blend from the microlensing model to the brightness of an MS star at an estimated distance from the lens. We summed the number of solutions where the brightness of the blend was smaller than the MS brightness and divided this number by the number of all solutions. The resulting number was interpreted as the probability that the lens is dark and not an MS star.
All input parameters are available in machine-readable form in an online archive attached to this paper.
@@ 7 Discussion and conclusions
We found in total 11 lenses for which the probability for the dark lens scenario for at least one solution exceeded 80% when we looked only at the Kroupa (2001) mass function. Eight of them passed all of the criteria imported in Section 4 (GaiaDR3-ULENS-025, GaiaDR3-ULENS-035, GaiaDR3-ULENS-069, GaiaDR3-ULENS-073, GaiaDR3-ULENS-088, GaiaDR3-ULENS-155, GaiaDR3-ULENS-343, and GaiaDR3-ULENS-353). Among the solutions that didn’t pass the $\pi_{\rm E}$ criterion, we found three more candidates (GaiaDR3-ULENS-103, GaiaDR3-ULENS-212, GaiaDR3-ULENS-331, and GaiaDR3-ULENS-155, but we accounted for this event in the first group). All but one of these events have Galactic coordinates towards the Galactic Centre. The estimated distance for 16 solutions (eight events) would suggest that they are located in the Galactic disc, rather than the Galactic bulge. For three solutions (two events), the estimated lens distance suggests a Galactic bulge lens. Three events seem to have the lens located within one kiloparsec from Earth. One event located towards the Galactic disc seems to have a lens no closer than 3.1 kpc, which means it would belong to the Scutum-Centaurus Arm of the Milky Way.
If we assumed that all of those 11 events are dark lenses, and instead follow a Mróz et al. (2021) mass function, we would end up with the majority of events with masses in the range of mass-gap objects and black holes. One event, GaiaDR3-ULENS-035, has mass consistent with a white dwarf, and three (GaiaDR3-ULENS-025, GaiaDR3-ULENS-155, GaiaDR3-ULENS-212) are consistent with a neutron star. Three events, GaiaDR3-ULENS-069, GaiaDR3-ULENS-103 and GaiaDR3-ULENS-343 overlap with the first mass gap in the one-sigma range. Four objects have larger masses (GaiaDR3-ULENS-073, GaiaDR3-ULENS-088, GaiaDR3-ULENS-331, GaiaDR3-ULENS-353) are consistent with BHs. All BH candidates have large uncertainties. There are issues with mass estimates for two of them: for GaiaDR3-ULENS-073 we struggled with finding the correct position of the RGC, which could result in the wrong source distance, leading to the wrong lens distance. GaiaDR3-ULENS-331 belongs to the group which didn’t pass the $\pi_{\rm E}$ criterion in Section 4.
@@
In the case of events where Gaia reported a proper motion, we calculated the transverse velocity. We present results for selected nine events in Figure 4. There is one high-velocity events: GaiaDR3-ULENS-212. This event has a relatively short Einstein time scale of 56 days. The estimated mass is consistent with an NS. The proper motion found in Gaia DR3 is not out of the ordinary, so this means that the lens would have to move fast to justify the relative proper motion and the resulting Einstein time scale.
All nine objects seem to separate into two groups: high-velocity, similar to the NS velocity from Hobbs et al. (2005), and low-velocity, similar to the velocity of a solitary BH from Lam & Lu (2023). The velocity doesn’t seem to depend on the lens mass, but the error bars are large and it is hard to draw conclusive statements. Moreover, here we are most likely getting an estimate of a velocity coming from a Galactic prior. More accurate estimates will be possible once we obtain an astrometric time series for these events.
Here we also measure only the most likely mass of candidate events. We cannot, however, confirm their nature as WDs, NS or BHs until we perform additional observations in other ranges of electromagnetic radiation, especially X-rays and UV. They could be some unusual objects, such as quark stars, or products of primordial black hole (PBH) mergers with other objects. Abramowicz et al. (2018) and Abramowicz et al. (2022) show a mechanism, that could produce a low-mass BH from a moon-mass PBH ($10^{25}\,\mathrm{g}>M_{\mathrm{PBH}}>10^{17}\,\mathrm{g}$) collision with an NS. Such low-mass BH would be in the mass range of an NS.
It is worth noting, that these are only candidates for dark lenses, and their mass measurement will be possible only when we will include astrometric time series. This data will be available only with Gaia DR4, no sooner than the end of 2025. This is an exciting prospect, as Gaia may allow us to observe previously unseen stellar populations. This becomes even more promising with the approaching start of the Vera C. Rubin Observatory and its Legacy Survey of Space and Time (Ivezić et al., 2019), as well as the launch of the Roman mission (Spergel et al., 2015; Akeson et al., 2019). Rubin will allow us to select long-duration microlensing events from the entire Galactic plane, while Roman can provide us with high-cadence astrometric and photometric observations in the Galactic bulge.
@@ Acknowledgements.
KK would like to thank Monika Sitek, Drs. Etienne Bachelet, Mariusz Gromadzki, Przemek Mróz, Radek Poleski, Milena Ratajczak, Rachel Street, Paweł Zieliński, Przemysław Mikołajczyk, as well as Profs. Michał Bejger, Wojciech Hellwing, Katarzyna Małek, and Łukasz Stawarz. This work was supported from the Polish NCN grants: Harmonia No. 2018/30/M/ST9/00311, Daina No. 2017/27/L/ST9/03221, and NCBiR grant within POWER program nr POWR.03.02.00-00-l001/16-00. LW acknowledges MNiSW grant DIR/WK/2018/12 and funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No. 101004719 (OPTICON-RadioNET Pilot, ORP). The MOA project is supported by JSPS KAKENHI Grant Number JP24253004, JP26247023, JP16H06287 and JP22H00153. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement. We acknowledge ESA Gaia, DPAC and the Photometric Science Alerts Team (http://gsaweb.ast.cam.ac.uk/alerts). This research has made use of publicly available data (https://kmtnet.kasi.re.kr/ulens/) from the KMTNet system operated by the Korea Astronomy and Space Science Institute (KASI) at three host sites of CTIO in Chile, SAAO in South Africa, and SSO in Australia. Data transfer from the host site to KASI was supported by the Korea Research Environment Open NETwork (KREONET).
@@ Appendix A Additional tables
In this appendix, we provide five tables:
@@
  • Table LABEL:tab:fit_data, which is the list of data sets used to obtain the final models for each of the 35 analysed events;
  • Table LABEL:tab:source_prop, which contains information about the source star;
  • Table LABEL:tab:lens_mass with the lens mass and distance estimates of the 14 candidate dark lens microlensing events;
  • Table LABEL:tab:lens_mass_edge with the lens mass and distance estimates of the five candidate dark lens microlensing events, that didn’t pass the $\pi_{\rm E}$ criterion, but were chosen to be analysed further.
  • Table LABEL:tab:fit_best_model, which is the list of parameters of best-fitting solutions of the 35 analysed events. This paper only provides the baseline magnitude and blending parameter for the $G$-band. A full, machine-readable version of this table, with all parameters, is available online.
Table 1: Table of the results of cross-match between 204 analysed Gaia DR3 events with other surveys. A full, machine-readable
version of this table is available at the CDS.
@@
{ThreePartTable}{TableNotes}
Notes: The table contains the event name, the name of the survey and the filter, the number of points used to obtain the best models, and the scale factor $k$ used to re-scale the photometric uncertainties. Light curves not used in the modelling process are marked with a strike-through. Light curves that were cleaned using the procedure outlined in Section 4 are marked in bold.
Table 2: The list of data sets used to obtain the final models for each of the 35 analysed events. The description of the columns is at the end of the table.
GaiaDR3- Survey, Num. $k$ GaiaDR3- Survey, Num. $k$
ULENS- filter of points ULENS- filter of points
003 018 Gaia, $G$ 24 1.00
Gaia, $G_{\rm BP}$ 22 1.00
Gaia, $G$ 27 1.00 Gaia, $G_{\rm RP}$ 22 3.13
Gaia, $G_{\rm BP}$ 27 1.00 OGLE, I 2780 1.02
Gaia, $G_{\rm RP}$ 27 1.00 KMTNet SAAO, I 563 4.47
ASASSN, V 84 1.79 KMTNet CTIO, I 840 1.97
KMTNet SSO, I 750 2.11
023 Gaia, $G$ 64 1.00 025 Gaia, $G$ 33
Gaia, $G_{\rm BP}$ 62 1.00 Gaia, $G_{\rm BP}$ 32
Gaia, $G_{\rm RP}$ 60 1.00 Gaia, $G_{\rm RP}$ 32
OGLE, I 127 1.50 OGLE, I 99
ASASSN, V 435 1.00
032 035 Gaia, $G$ 36 1.00
Gaia, $G$ 40 1.00 Gaia, $G_{\rm BP}$ 33 1.00
Gaia, $G_{\rm BP}$ 38 1.00 Gaia, $G_{\rm RP}$ 33 1.00
Gaia, $G_{\rm RP}$ 37 1.00 OGLE, I 1063 1.00
ASASSN, V 308 2.09 MOA, I 19531 1.63
MOA, V 612 1.89
047 Gaia, $G$ 37 1.00 57
Gaia, $G_{\rm BP}$ 34 1.00 Gaia, $G$ 24
Gaia, $G_{\rm RP}$ 35 1.00 Gaia, $G_{\rm BP}$ 24
KMTNet SAAO, I 2051 6.69 Gaia, $G_{\rm BP}$ 23
KMTNet CTIO, I 3871 1.00
KMTNet SSO, I 2726 4.77
069 Gaia, $G$ 36 1.00 073 Gaia, $G$ 22 1.00
Gaia, $G_{\rm BP}$ 31 1.00 Gaia, $G_{\rm BP}$ 20
Gaia, $G_{\rm RP}$ 32 1.00 Gaia, $G_{\rm RP}$ 19 1.49
KMTNet SAAO, I 549 OGLE, I 368 1.01
KMTNet CTIO, I 820 1.81 MOA, I 2068 1.71
KMTNet SSO, I 736 4.22 MOA, V 277 1.57
MOA, I 2068 1.83
MOA, V 277 2.64
078 079 Gaia, $G$ 24 5.41
Gaia, $G_{\rm BP}$ 25 1.59
Gaia, $G$ 39 1.00 Gaia, $G_{\rm RP}$ 24 1.47
Gaia, $G_{\rm BP}$ 38 1.00 OGLE, I 600 1.19
Gaia, $G_{\rm RP}$ 38 1.76 KMTNet SAAO, I 2923 4.59
OGLE, I 730 1.01 KMTNet CTIO, I 4327 2.01
KMTNet SSO, I 2678 2.97
MOA, I 4315 1.57
MOA, V 334 3.00
088 Gaia, $G$ 33 1.00 089
Gaia, $G_{\rm BP}$ 32 Gaia, $G$ 24 1.00
Gaia, $G_{\rm RP}$ 33 1.00 Gaia, $G_{\rm BP}$ 25
OGLE, I 682 1.01 Gaia, $G_{\rm RP}$ 23 1.00
KMTNet SAAO, I 548 3.53 OGLE, I 934 1.00
KMTNet CTIO, I 837 3.36 KMTNet SAAO, I 242 2.88
KMTNet SSO, I 748 3.45 KMTNet CTIO, I 385 1.87
MOA, I 7974 1.36 KMTNet SSO, I 341 2.53
MOA, V 281 1.42
097 Gaia, $G$ 54 1.00 103 Gaia, $G$ 66 1.00
Gaia, $G_{\rm BP}$ 52 Gaia, $G_{\rm BP}$ 59 1.00
Gaia, $G_{\rm RP}$ 52 1.00 Gaia, $G_{\rm RP}$ 63 1.00
OGLE, I 791 1.00 OGLE, I 144 1.19
118 127 Gaia, $G$ 27 1.16
Gaia, $G$ 22 1.00 Gaia, $G_{\rm BP}$ 16
Gaia, $G_{\rm BP}$ 22 1.00 Gaia, $G_{\rm RP}$ 16
Gaia, $G_{\rm RP}$ 23 OGLE, I 2677 1.01
OGLE, I 168 1.30 KMTNet SAAO, I 288 3.41
ASASSN, V 23 2.55 KMTNet CTIO, I 409 3.83
KMTNet SSO, I 377 1.71
142 Gaia, $G$ 29 1.00 143 Gaia, $G$ 60
Gaia, $G_{\rm BP}$ 27 Gaia, $G_{\rm BP}$ 58
Gaia, $G_{\rm RP}$ 25 1.00 Gaia, $G$ 60
OGLE, I 638 1.44 OGLE, I 122
155 Gaia, $G$ 47 1.00 196
Gaia, $G_{\rm BP}$ 45
Gaia, $G_{\rm RP}$ 45 1.00 Gaia, $G$ 31 1.00
OGLE, I 329 1.00 Gaia, $G_{\rm BP}$ 25
KMTNet SAAO, I 237 5.16 Gaia, $G_{\rm RP}$ 31 1.00
KMTNet CTIO, I 373 2.08 OGLE, I 956 1.06
KMTNet SSO, I 329 3.56
MOA, I 1881 1.71
MOA, V 237 1.22
212 Gaia, $G$ 59 1.00 230
Gaia, $G_{\rm BP}$ 54 1.00
Gaia, $G_{\rm RP}$ 55 1.25
GSA, $G$ 80 1.00 Gaia, $G$ 28 1.00
OGLE, I 2140 1.21 Gaia, $G_{\rm BP}$ 24
KMTNet SAAO, I 694 5.16 Gaia, $G_{\rm RP}$ 25
KMTNet CTIO, I 956 2.37 OGLE, I 2838 1.20
KMTNet SSO, I 855 1.97
259 270 Gaia, $G$ 24 1.00
Gaia, $G$ 58 1.00 Gaia, $G_{\rm BP}$ 9
Gaia, $G_{\rm BP}$ 50 Gaia, $G_{\rm RP}$ 9
Gaia, $G_{\rm RP}$ 50 1.00 OGLE, I 1140 1.00
GSA, $G$ 164 1.00 KMTNet SAAO, I 551 1.91
OGLE, I 191 1.20 KMTNet CTIO, I 820 1.27
KMTNet SSO, I 759 1.45
275 284 Gaia, $G$ 42 1.00
Gaia, $G$ 40 1.00 Gaia, $G_{\rm BP}$ 34
Gaia, $G_{\rm BP}$ 38 Gaia, $G_{\rm RP}$ 37
Gaia, $G_{\rm RP}$ 39 1.00 OGLE, I 1095 1.00
OGLE, I 131 1.24 KMTNet SAAO, I 748 3.68
KMTNet CTIO, I 992 2.24
KMTNet SSO, I 921 2.69
326 Gaia, $G$ 39 1.00 331 Gaia, $G$ 42 1.00
Gaia, $G_{\rm BP}$ 22 Gaia, $G_{\rm BP}$ 22
Gaia, $G_{\rm RP}$ 33 Gaia, $G_{\rm RP}$ 36 1.00
OGLE, I 791 1.00 OGLE, I 1776 1.06
KMTNet SAAO, I 531 2.00 KMTNet SAAO, I 527 1.91
KMTNet CTIO, I 823 1.40 KMTNet CTIO, I 822 1.20
KMTNet SSO, I 731 1.34 KMTNet SSO, I 755 1.44
343 Gaia, $G$ 20 1.00 353 Gaia, $G$ 18 1.00
Gaia, $G_{\rm BP}$ 10 Gaia, $G_{\rm BP}$ 12
Gaia, $G_{\rm RP}$ 10 Gaia, $G_{\rm RP}$ 12
OGLE, I 725 1.00 OGLE, I 506 1.00
KMTNet SAAO, I 679 2.50 KMTNet SAAO, I 657 2.22
KMTNet CTIO, I 919 1.81 KMTNet CTIO, I 241
KMTNet SSO, I 841 1.79 KMTNet SSO, I 841 1.50
MOA, I 10498 1.21 MOA, I 5223 1.13
MOA, V 328 1.15 MOA, V 191 1.14
359 Gaia, $G$ 21 1.00 363 Gaia, $G$ 23 1.00
Gaia, $G_{\rm BP}$ 17 Gaia, $G_{\rm BP}$ 19
Gaia, $G_{\rm RP}$ 19 1.00 Gaia, $G_{\rm RP}$ 22
OGLE, I 553 1.00 OGLE, I 824 1.00
KMTNet SAAO, I 260 2.79 KMTNet SAAO, I 780 1.68
KMTNet CTIO, I 994 1.58 KMTNet CTIO, I 386 1.39
KMTNet SSO, I 895 1.88 KMTNet SSO, I 364 1.45
[…]